# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ 3.4 Vector-borne disease simulation results
The Aedes aegypti mosquito typically flies up to 400 meters in search of water-filled containers to lay eggs, but it generally stays near human habitation [38]. Our simulation used a 500x500 square meter patch size to represent the mosquito’s activity area. The other simulation parameters or ranges are specified in Table. 4.
3.4.1 Disease progression in an uncontrolled environment
@@
Figure 24: (a) Infectious disease states of non-hospitalized agents, (b) infectious disease states of hospitalized agents, (c) infectious disease states of all agents, and (d) major states changes of all agents over time.
At the start of the simulation, changes in the disease state of agents can be observed over time, as shown in Fig. 24. Analyzing the rate of change indicates that once dengue starts spreading, the infection rate gradually increases before eventually declining. However, in this simulation, each individual can only be infected once. If reinfection were allowed, the infection rates would likely be higher. Reinfection could be modeled by resetting recovered agents to a susceptible state, but for simplicity, this is not considered in this simulation.
3.4.2 Disease progression in a controlled environment
Various strategies can be employed to control dengue. In this simulation, the focus is on managing mosquito populations. If the cumulative number of exposed agents in a zone over an $n$-day period exceeds a predefined threshold, $E_{h}^{k}$, the vector population in the affected zone’s patches is reduced by $m\%$. This reduction represents interventions such as mosquito spraying and the elimination of breeding sites. Fig. 25 illustrates a case where $n=7$, $E_{h}^{k}=2$, and $m=75$.
Figure 25: The infected vector count changes over 2 months in (a) an uncontrolled and (b) controlled environment.
Fig. 26 illustrates the mosquito control strategy under different scenarios. As the infection count increases, the number of exposed individuals in each zone also rises. At the end of each week, zones with more than n exposed agents undergo vector reduction. However, when higher thresholds for exposed counts are allowed, it becomes evident that this strategy is not significantly different from the uncontrolled case.
@@ 3.4.2 Disease progression in a controlled environment
Figure 26: The effect of controlling mosquitoes
3.4.3 Effect of temperature
As mentioned in Section 2.9.1, the average rate of progression of vectors from the exposed state to the infectious state depends on the mosquito’s incubation time. The mosquito incubation time is highly influenced by temperature. Fig. 27 shows how the total infectious agent count changes as the mosquito incubation period varies with temperature. During high-temperature seasons, it is important to implement necessary measures to reduce the propagation of dengue.
Figure 27: Effect of temperature on dengue disease progression
Funding
This research was funded by the International Development Research Centre, Canada, grant number 109586-001.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
4 Conclusion
The emerging threat of air-borne and vector-borne diseases has existed forever in human history. Recent events such as the COVID-19 pandemic and the continuous spread of dengue fever are key examples of this. Some countries have faced severe consequences since they were not able to manage the spread of the disease by executing effective policies. In our study, we present AVSim, an agent-based model designed to assist policymakers in making more informed decisions for the effective control of diseases by taking in to account numerous possible scenarios.
AVSim sets itself apart by its ability to model complex real-world environments, a feature many other agent-based models lack. It can precisely simulate entire cities, subdivisions, zones, and even individual households. By leveraging demographic data of the modeled geographical area, AVSim generates agent entities that accurately represent the population.
The movement of these agents is modeled using a novel probability-based approach informed by real-world GPS data collected from 100 individuals across various occupational classes. This innovative method effectively replicates human behavior and has been rigorously validated through an extensive study of human motion patterns.
Additionally, AVSim incorporates a realistic transport network, enabling agents to move via modeled transportation systems. It can simulate infections based on professional classes and supports a variety of analyses to study disease dynamics. Together, these capabilities place AVSim in a class of its own, far surpassing the limitations of other agent-based models.
The results from AVSim enable users to identify vulnerable agent groups and high-risk locations. Preventive measures, such as promoting safe hygiene practices and implementing vaccination campaigns, can then be simulated to assess their impact on mitigating disease spread. These simulations provide valuable insights, helping policymakers make informed decisions and implement strategies to control disease transmission while minimizing societal and economic disruption in the region.
[…]

### v2

@@ IV-D Vector-borne disease simulation results
Our simulation used a 500x500 square meter patch size to represent the mosquito’s activity area, as Aedes aegypti mosquitoes typically fly up to 400 meters [35]. The other simulation parameters are set as follows: $\psi_{v}=0.3$, $\mu_{v}=1/14$, $\sigma_{v}=0.5$, $\sigma_{h}=10$, $K_{v}\in[100,200]$, $\beta_{vh}=0.33$, and $\beta_{hv}=0.33$.
Figure 13: Infectious disease states of (a) non-hospitalized, (b) hospitalized (only symptomatic agents are hospitalized), (c) all agents, and (d) major disease states of all agents over time.
IV-D1 Disease progression in an uncontrolled environment
@@
As the simulation progresses, changes in the disease state of agents can be observed over time, as shown in Fig. 13. Analyzing the rate of change indicates that once dengue starts spreading, the infection rate gradually increases before eventually declining. However, in this simulation, each individual can only be infected once. If reinfection were allowed, the infection rates would likely be higher. Reinfection could be modeled by resetting recovered agents to a susceptible state, but for simplicity, this is not considered in this simulation.
IV-D2 Disease progression in a controlled environment
Various strategies can be employed to control dengue. In this simulation, the focus is on managing mosquito populations. If the cumulative number of exposed agents in a zone over an $n$-day period exceeds a predefined threshold, $E_{h}^{k}$, the vector population in the affected zone’s patches is reduced by $m\%$. This reduction represents interventions such as mosquito spraying and the elimination of breeding sites. Fig. 14 illustrates a comparison between uncontrolled spread and a control mechanism where $n=7$, $E_{h}^{k}=2$, and $m=75$.
Figure 14: The infected vector count changes over 2 months in (a) a controlled and (b) an uncontrolled environment.
Figure 15: The sum of Infectious, Mild, Severe, and Asymptomatic cases changes with a) the control strategies and b) temperature
Fig. 15a illustrates the mosquito control strategy under different scenarios. As the infection count increases, the number of exposed individuals in each zone also rises. At the end of each week, zones with more than n exposed agents undergo vector reduction. However, when higher thresholds for exposed counts are allowed, it becomes evident that this strategy is not significantly different from the uncontrolled case.
@@ IV-D2 Disease progression in a controlled environment
IV-D3 Effect of temperature
As mentioned in Section III-H1, the average rate of progression of vectors from the exposed state to the infectious state depends on the mosquito’s incubation time. The mosquito incubation time is highly influenced by temperature. Fig. 15b shows how the total infectious agent count changes as the mosquito incubation period varies with temperature. During high-temperature seasons, it is important to implement necessary measures to reduce the propagation of dengue.
Similarly, environmental factors such as rainfall also play a significant role in dengue transmission. During the rainy season, increased mosquito emergence rates and higher carrying capacities contribute to a surge in dengue cases. Policymakers can analyze historical data and use simulations to predict dengue outbreaks in advance, allowing them to take necessary preventive measures for the future.
V Conclusion
Airborne and vector-borne diseases, such as COVID-19 and dengue fever, pose persistent threats to humanity. Recent pandemics have exposed the severe consequences of inadequate disease control policies. Effective containment strategies require more than generalized models that rely on simulations grounded in how people move, interact, and respond across different social and spatial contexts.
AVSim addresses this need by embedding real-world mobility and behavioral patterns into an agent-based framework capable of simulating disease dynamics with occupational and environmental specificity. By grounding agent behavior in empirical GPS data and uncovering latent structure in daily routines, AVSim provides a high-resolution lens into how infections propagate through human and vector-mediated contact. Its ability to represent both airborne and vector-borne disease mechanisms while accounting for geography, transport networks, and heterogeneous behaviors enables a more nuanced assessment of transmission pathways and intervention outcomes.
Through simulations of COVID-19 and dengue, AVSim demonstrates how grounded behavioral modeling can reveal vulnerabilities in population structure, identify spatial hotspots, and evaluate the effectiveness of targeted interventions such as quarantine, vaccination, and vector control. As public health challenges become increasingly complex, AVSim provides a scalable and adaptable platform for exploring disease dynamics in a manner that is both realistic and actionable.